"In 3.5 h, a balloon drifts 21.5 km north, 9.70 km east, and 2.88 km upward from its release point on the ground. Find (a) the magnitude of its average velocity and (b) the angle its average velocity m

Grade:12th pass

1 Answers

Vikas TU
7 years ago
First Part =>
Avg. Velocity = Total Displacement / Total Time 
Net displc. => root(21.5^2 + 9.70^2) = >root(462.5 +  94.09) => root(556.59) = > 23.59 Km.
Now also, with upward distance => root(2.88^2 + 23.59^2) = > root(556.5 + 8.29) => root(556.7944) =>  23.77 km.
 
= 23.77/3.5 = > 6.79 km/hr.
 
Angle => thetha = arccos(2.88/23.59) => arccos(0.122) = > 1.4488 degree.